Consider the equation

1/u + 1/v = 1/f

Given that the measured values of u and v are u = 0.29 ± 0.02 and v = 0.65 ± 0.03 calculate the greatest possible error in estimating f.
Answer #1

Ý + t = ½ u = 0.2.97 +0.02 V = 0.65 ± 0.03 = 4 + + 0.1885 0.99 + utv (0-0230.03) + = + > 0.02 1ro- 29 + 0.03 0.65 » f = uv (0

Answer is in the rectangular box at the end